OpenStudy (anonymous):

xy = −25. What is the constant of variation for this relationship? A. 1/4 B. -25 C. -2.5 D. -0.4

OpenStudy (anonymous):

@tester97

OpenStudy (tester97):

i dont remember constant variation sorry >.<

OpenStudy (anonymous):

V varies inversely with T and V = 14 when T = 7. Which equation shows this relationship?

OpenStudy (anonymous):

can u help me with that one? ^^

OpenStudy (tester97):

its the same as the last one we did just with different letters

OpenStudy (tester97):

K T=------- V

OpenStudy (anonymous):

A. v = 2/T B. v = 98/T C. v = 2T D. v = 98T

OpenStudy (tester97):

c